For a series LRC circuit, if L=30H, R= 0.05 ohms and C= 5.8F, what is the time required for the amplitude of the charge stored by the capacitor to be halved?


1
Expert's answer
2021-11-15T17:51:09-0500

We know that

Charge

"q=q_0(1-e^{\frac{-t}{RC}})"

"q=q_0(1-e^{\frac{-t}{0.05\times5.8}})"

"\frac{q_0}{2}=q_0(1-e^{\frac{-t}{0.05\times5.8}})"

"\frac{1}{2}=e^{-\frac{t}{0.05\times5.8}}"

"ln2=\frac{t}{0.05\times5.8}"

"t=0.29\times ln2=0.202sec"